# Thompson Sampling for Repeated Newsvendor

### v1

@@ 2.2. Preliminaries
Demand Distribution: Newsvendor Family. The newsvendor (newsboy) family, introduced by (Braden and Freimer, 1991), is known to be the only family whose posterior distributions with censored demand information have conjugate priors. Formally put, a random variable is a member of the newsvendor distributions if its density is given by
$f_{\theta}(x)=\theta d^{\prime}(x)e^{-\theta d(x)},\quad F_{\theta}(x)=1-e^{-\theta d(x)},$
where $d^{\prime}(x)>0,\ \forall x>0$, so $f_{\theta}(x)$ is positive on $(0,\infty)$ $\lim_{x\rightarrow 0}d(x)=0$ and $\lim_{x\rightarrow\infty}d(x)=\infty$. So $F_{\theta}(x)$ is a valid probability distribution, where $d(x)$ is a positive, differentiable, and increasing function and $\theta\in R_{+}$.
(Lariviere and Porteus, 1999) show that when the demand distribution is Weibull with a gamma prior, the optimal solution for repeated newsvendor problem admits a closed form. Namely, by letting $d(x)=x^{k}$ with a known constant $k>0$, we get the Weibull distribution. If $k=1$, we get the exponential distribution. In such cases, the underlying density function of demand is
$f_{\theta}(x)=\theta kx^{k-1}e^{-\theta x^{k}}.$
Prior Distribution and Parametric Demand. With the true value of $\theta_{\star}$ being unknown, the decision maker initiates TS with a prior distribution $\rho_{0}$ at the outset. Throughout the paper, we adopt the prior family and parametric demand introduced by (Braden and Freimer, 1991). Namely, the prior follows $\rho_{0}\sim\operatorname{Gamma}(\alpha_{0},\beta_{0})$ ($\rho_{0}(\theta)=\frac{\beta_{0}^{\alpha_{0}}}{\Gamma(\alpha_{0})}\theta^{\alpha_{0}-1}e^{-\theta\beta_{0}})$. When demand is described by a member of the newsvendor family, the gamma distribution remains a conjugate prior. Under the Weibull distribution of demand, we have
$F^{-1}_{\theta_{\star}}(\frac{p}{p+h})=\frac{1}{\theta_{\star}}\left(-\ln(\frac{h}{p+h})\right)^{1/k},$
where $F^{-1}_{\theta_{\star}}$ is the inverse cumulative distribution function. Here, we emphasize that this prior is only used to initiate TS and we do not impose any prior distribution on $\theta_{*}$.
@@
Consider we have $t\geq 2$ observations of data pairs denoted as $Y=(Y_{0},Y_{1},\cdots,Y_{t}),\delta=(\delta_{0},\cdots,\delta_{t})$ and we use $C$ denote the set of all observations of censored data pairs and $\bar{C}$ denote the set of all observations of uncensored data pairs , where $|C|=m$, $|\bar{C}|=n$, and $m+n=t$ Then the likelihood function is
$\mathcal{L}\left(\theta\mid Y,\delta\right)$ $=\Pi_{i=1}^{n}\left(f_{\theta}\left(Y_{i}\right)\right)^{i}\Pi_{j=1}^{m}\left((1-F_{\theta}\left(Y_{j}\right)\right)^{j}$
$=\left(\theta k\right)^{n}\left(\Pi_{i=1}^{n}Y_{i}\right)^{k-1}e^{-\theta\sum_{l=1}^{t}Y_{l}^{k}}.$
Posterior Update. The posterior demand distribution $\rho_{t}$ at the beginning of period $t$ can be derived as follows:
@@
$\propto\beta_{0}^{\alpha_{0}}\theta^{\alpha_{0}-1}e^{-\theta\beta_{0}}\times\left(\theta k\right)^{n}\left(\Pi_{i=1}^{n}Y_{i}\right)^{k-1}e^{-\theta\sum_{l=1}^{t}Y_{l}^{k}}$
$\propto\theta^{\alpha_{0}+n-1}e^{-\theta\left(\beta_{0}+\sum_{i=1}^{t}Y_{l}^{k}\right)}$
@@
2.3. Algorithm: Thompson Sampling for Repeated Newsvendor Problem
TS is a Bayesian approach used to balance exploration and exploitation in sequential decision-making problems. In the context of the newsvendor problem, TS can be implemented to decide on the optimal order quantity under demand uncertainty. The specific TS procedure involves the following steps:
@@ 2.3. Algorithm: Thompson Sampling for Repeated Newsvendor Problem
       Place order quantity
$y_{t}=\frac{1}{\theta_{t}}\left(-\ln\left(\frac{h}{p+h}\right)\right)^{1/k},$
where $\theta_{t}\sim\operatorname{Gamma}(\alpha_{t},\beta_{t})$;
       Observe sales $Y_{t}=\min\{D_{t},y_{t}\}$ and indicator of whether demand is censored $\delta_{t}=\mathbf{1}[D_{t}<y_{t}]$ ;
       Update the posterior $\rho_{t}\sim\operatorname{Gamma}\left(\alpha_{t},\beta_{t}\right)$, where
$\alpha_{t}=\alpha_{0}+\sum_{i=0}^{t-1}\delta_{i},\quad\beta_{t}=\beta_{0}+\sum_{i=0}^{t-1}Y_{i}^{k}.$
@@
Initially, the environment draws a sample of $\theta_{\star}$ from prior $\rho_{0}=\operatorname{Gamma}(\alpha_{0},\beta_{0})$, which is unknown to DM. and a known time horizon $T$. Then, for each $t\in[T]$, DM place the order quantity $y_{t}$ and then observes the sales $Y_{t}$, which is the minimum of demand and order quantity. Then the posterior is updated accordingly. $y_{t}$ iteratively updates the posterior and samples from it. Specifically, $y_{t}(\theta_{t})=F^{-1}_{\theta_{t}}(\frac{p}{p+h})=\frac{1}{\theta_{t}}\left(-\ln\left(\frac{h}{p+h}\right)\right)^{1/k}$ and the property of $\theta_{t}$ (e.g. $\mathbb{E}[\theta_{t}]$ and $\mathbb{E}[1/\theta_{t}]$). $\theta_{t}$ is sampled from Gamma distribution with $\alpha_{t}$, $\beta_{t}$. This is motivated from the posterior update. $\theta_{t}$ satisfies $\operatorname{Gamma}(\alpha_{t},\beta_{t})$ $\mathbb{E}[1/\theta_{t}]=\frac{\beta_{t}}{\alpha_{t}-1}$. TS efficiently balances exploration (learning about the true demand distribution) and exploitation (placing optimal orders based on current knowledge). This approach is particularly useful in multi-period inventory problems, where demand is uncertain and needs to be learned over time.
3. Regret Analysis
In this section, we provide the analysis for the regret upper bound on our Algorithm 1, which is equal to
$\tilde{O}\left(\max\{h,p\}\cdot\left(-\ln(\frac{h}{p+h})\right)^{\frac{1}{k}}\cdot\frac{1}{\theta_{\star}^{2}}\cdot\sqrt{T}\right).$
Section 3.1 we provide the main theorem that state the upper bound. In Section 3.2 we provide a sketch proof for proving the Theorem. The proof consists of three key steps:
  • Lipchitz Continuity of Regret (Section 3.2.1)
  • Confidence Analysis of Estimation (Section 3.2.2)
  • Lower bounding the Actions (Section 3.2.3)
We also discuss how the these steps can be generalized to broader models of online learning with censored feedback in Section 5.
3.1. Main Result: Regret
Theorem 3.1.
@@ 3.1. Main Result: Regret
3.2. Proof for Theorem 3.1
The entire proof consists of several main steps. Firstly, we focus on $\operatorname{Regret(T,\theta_{\star})}$ for a fixed $\theta_{*}$. We decompose it by Lipchitz Continuity.
3.2.1. Key Step 1: Regret Decomposition: Lipchitz Continuity

### v2

@@ 2.2 Preliminaries
Demand Distribution: Newsvendor Family. The newsvendor (newsboy) family, introduced by Braden and Freimer (1991), is known to be the only family whose posterior distributions with censored demand information have conjugate priors. A random variable is a member of the newsvendor distributions if its density is given by $f_{\theta}(x)=\theta d^{\prime}(x)e^{-\theta d(x)},\quad F_{\theta}(x)=1-e^{-\theta d(x)},$ where $d^{\prime}(x)>0,\ \forall x>0$, so $f_{\theta}(x)$ is positive on $(0,\infty)$ $\lim_{x\rightarrow 0}d(x)=0$ and $\lim_{x\rightarrow\infty}d(x)=\infty$. So $F_{\theta}(x)$ is a valid probability distribution, where $d(x)$ is a positive, differentiable, and increasing function and $\theta\in R_{+}$. Lariviere and Porteus (1999) show that when the demand distribution is Weibull with a gamma prior, the optimal solution for repeated newsvendor problem admits a closed form.
Prior Distribution and Parametric Demand. With the true value of $\theta_{\star}$ being unknown, the decision maker initiates TS with a prior distribution $\rho_{0}$ at the outset. Throughout the paper, we adopt the prior family and parametric demand introduced by Braden and Freimer (1991). Namely, the prior follows $\rho_{0}\sim\operatorname{Gamma}(\alpha_{0},\beta_{0})$ ($\rho_{0}(\theta)=\frac{\beta_{0}^{\alpha_{0}}}{\Gamma(\alpha_{0})}\theta^{\alpha_{0}-1}e^{-\theta\beta_{0}})$. When demand is described by a member of the newsvendor family, the gamma distribution remains a conjugate prior. Under the Weibull distribution of demand, we have $F^{-1}_{\theta_{\star}}(\frac{p}{p+h})=\left(\frac{1}{\theta_{t}}\right)^{1/k}\left(-\ln(\frac{h}{p+h})\right)^{1/k},$ where $F^{-1}_{\theta_{\star}}$ is the inverse cumulative distribution function. Here, we emphasize that this prior is only used to initiate TS and we do not impose any prior distribution on $\theta_{\star}$.
@@
Consider we have $t\in[T]$ observations of data pairs denoted as $Y=(Y_{0},Y_{1},\cdots,Y_{t}),\delta=(\delta_{0},\cdots,\delta_{t})$ and we use $C$ denote the set of all observations of censored data pairs and $\bar{C}$ denote the set of all observations of uncensored data pairs , where $|C|=m$, $|\bar{C}|=n$, and $m+n=t$. Then the likelihood function is
$\mathcal{L}\left(\theta\mid Y,\delta\right)=\Pi_{i=1}^{n}\left(f_{\theta}\left(Y_{i}\right)\right)^{i}\Pi_{j=1}^{m}\left((1-F_{\theta}\left(Y_{j}\right)\right)^{j}=\left(\theta k\right)^{n}\left(\Pi_{i=1}^{n}Y_{i}\right)^{k-1}e^{-\theta\sum_{l=1}^{t}Y_{l}^{k}}.$
Posterior Update. The posterior demand distribution $\rho_{t}$ at the beginning of period $t$ can be derived as:
@@
2.3 Algorithm: Thompson Sampling for Repeated Newsvendor Problem
The TS algorithm is presented in 1. TS is a Bayesian approach used to balance exploration and exploitation in sequential decision-making problems. In the context of the newsvendor problem, TS can be implemented to decide on the optimal order quantity under demand uncertainty.
@@ 2.3 Algorithm: Thompson Sampling for Repeated Newsvendor Problem
   Place order quantity $y_{t}=\left(\frac{1}{\theta_{t}}\right)^{1/k}\left(-\ln(\frac{h}{p+h})\right)^{1/k},$ where $\theta_{t}\sim\operatorname{Gamma}(\alpha_{t-1},\beta_{t-1})$;
   Observe sales $Y_{t}=\min\{D_{t},y_{t}\}$ and indicator of whether demand is censored $\delta_{t}=\mathbf{1}[D_{t}<y_{t}];$ Update the posterior $\rho_{t}\sim\operatorname{Gamma}\left(\alpha_{t},\beta_{t}\right)$, where $\alpha_{t}=\alpha_{0}+\sum_{i=1}^{t}\delta_{i},\ \beta_{t}=\beta_{0}+\sum_{i=1}^{t}Y_{i}^{k}.$
@@
Initially, the environment draws a sample of $\theta_{\star}$ from prior $\rho_{0}=\operatorname{Gamma}(\alpha_{0},\beta_{0})$, which is unknown to retailer. and a known time horizon $T$. Then, for each $t\in[T]$, retailer places the order quantity $y_{t}$ and then observes the sales $Y_{t}$, which is the minimum of demand and order quantity. Then the posterior is updated accordingly. $y_{t}$ iteratively updates the posterior and samples from it. Specifically, $y_{t}=F^{-1}_{\theta_{t}}(\frac{p}{p+h})=\left(\frac{1}{\theta_{t}}\right)^{1/k}\left(-\ln(\frac{h}{p+h})\right)^{1/k}$ and the property of $\theta_{t}$. We sample $\theta_{t}$ from a Gamma distribution with shape parameter $\alpha_{t}$ and rate parameter $\beta_{t}$, i.e., $\theta_{t}\sim\text{Gamma}(\alpha_{t},\beta_{t})$. This choice is motivated by conjugate prior properties in the posterior update. $\mathbb{E}[\frac{1}{\theta_{t}}]=\frac{\beta_{t}}{\alpha_{t}-1}$. TS efficiently balances exploration (learning about the true demand distribution) and exploitation (placing optimal orders based on current knowledge).
3 Frequentist Regret Analysis for Weibull Distribution
In this section, we provide the analysis for the regret upper bound on our Algorithm 1, which is shown in the following theorem:
Theorem 1.
@@ 3 Frequentist Regret Analysis for Weibull Distribution
We provide a sketch proof for proving the Theorem 1. The proof consists of three key steps: 1. Lipchitz Continuity of Regret (Section 3.2.1), 2. Confidence Analysis of Estimation (Section 3.2), and 3. Lower bounding the Actions (Section 3.1). We also discuss how the these steps can be generalized to broader models of online learning with censored feedback in Section A.1.
3.1 Key Step 1: Uniform Lower Bound of $y_{t}$
In order to establish the regret bound, it is essential to first prove a uniform lower bound for $y_{t}$, as this will play a crucial role in subsequent derivations. According to Lemma 1, we have:
Lemma 1.
The proof of Lemma 1 is provided in Appendix A.4. Building upon this lemma, we proceed by conditioning on the event that $\frac{1}{\theta_{t}}>\frac{\beta_{t}}{2\alpha_{t}}$ and that the demand $D_{t}$ satisfies $D_{t}\geq\underline{D}$. Under these conditions, we can derive a lower bound for $y_{t}$ as follows:
$y_{t}$ $=\left(\frac{1}{\theta_{t}}\right)^{1/k}\left(-\ln(\frac{h}{p+h})\right)^{1/k}$
$\geq\left(\frac{\beta_{t}}{2\alpha_{t}}\right)^{1/k}\cdot\left(-\ln(\frac{h}{p+h})\right)^{1/k}$ (2a)
$=\left(\frac{1}{2}\right)^{1/k}\left(-\ln(\frac{h}{p+h})\right)^{1/k}\cdot\left(\frac{\beta_{0}+\sum_{i=1}^{t}\min\{y_{i},D_{i}\}^{k}}{\alpha_{0}+\sum_{i=1}^{t}\delta_{i}}\right)^{1/k}$ (2b)
$\geq\left(-\frac{1}{2}\ln(\frac{h}{p+h})\right)^{1/k}\cdot\min\left\{\left(\frac{\beta_{0}}{\alpha_{0}}\right)^{1/k},\underline{D}\right\}=L.$ (2c)
Here, equation (2a) follows directly from the definition of $y_{t}$ in equation (4) together with Lemma 1, which guarantees that $\left(1/\theta_{t}\right)^{1/k}$ is, with high probability, bounded below by $\left(\beta_{t}/(2\alpha_{t})\right)^{1/k}$. Equality (2b) then uses the update rules for $\alpha_{t}$ and $\beta_{t}$ in Algorithm 1. Finally, inequality (2c) is obtained by applying Lemma 2, which ensures the ratio is uniformly bounded from below, thereby establishing the constant lower bound $L$.
Lemma 2.
From the closed-form expression (2b) and Lemma 2, we reveal the most important ascept of TS algorithm as follows:
  1. 1.
  2. 2.
    when $\delta_{i}=0$ (i.e. $D_{t}\geq y_{t}$). we get the censored demand, which indicates the past action is relatively small. Interestingly, since $\delta_{i}$ appears in the denominator in the closed-form expression (2b), TS naturally pushes future actions higher in subsequent periods, preventing the algorithm from getting stuck with poor estimates.
Applying Lemma 2 ( proved in Appendix A.5 ) in our context, and considering that $D_{t}\geq\underline{D}$, we conclude that $y_{t}$ is uniformly bounded below by $L$ for all $t$.
This uniform lower bound on $y_{t}$ is a critical to establish the regret bound. we plug back the lower bound to the definition of $\alpha_{t}$ in Lemma 6 to analyze the term $\left|\alpha_{t}-1\right|$, which analysis is referred in Appendix A.9. From the closed-form expression (2c) and Lemma 2, we reveal the most important aspect of the TS algorithm as follows:
  1. 1.
    When $\delta_{i}=1$ (i.e. $D_{t}<y_{t}$), we obtain the full demand observation. The increment is uncensored and can provide accurate information about the demand’s upper tail.
  2. 2.
    When $\delta_{i}=0$ (i.e. $D_{t}\geq y_{t}$), we get censored demand, which indicates the past action was relatively small. Interestingly, since $\delta_{i}$ appears in the denominator in (2b), TS naturally pushes future actions higher in subsequent periods, preventing the algorithm from getting stuck with poor estimates.
This key observation illustrates how TS automatically balances the exploration-exploitation trade-off in the repeated newsvendor problem. Applying Lemma 2 (proved in Appendix A.5) in our context, and considering that $D_{t}\geq\underline{D}$, we conclude that $y_{t}$ is uniformly bounded below by $L$ for all $t$. This uniform lower bound on $y_{t}$ is critical for establishing the regret bound.
3.2 Key Step 2: Confidence Analysis of $\left|\mathbb{E}[y_{t}]-y_{\star}\right|$
Before proceeding, we give the definition for $y_{t}$ and $y_{\star}$:
Lemma 3.
The order quantity $y_{t}$ at $t$ and the optimal myopic order quantity $y_{\star}$ satisfy
$y_{\star}(\theta_{\star})$ $=F^{-1}_{\theta_{\star}}(\tfrac{p}{p+h})=\left(\frac{1}{\theta_{\star}}\right)^{1/k}\left(-\ln(\tfrac{h}{p+h})\right)^{1/k}$ (3)
$y_{t}(\theta_{t})$ $=F^{-1}_{\theta_{t}}(\tfrac{p}{p+h})=\left(\frac{1}{\theta_{t}}\right)^{1/k}\left(-\ln(\tfrac{h}{p+h})\right)^{1/k}$ (4)
where $F^{-1}$ is the inverse cumulative distribution function of the demand distribution. Moreover, $\mathbb{E}\left[\tfrac{1}{\theta_{t}}\right]=\tfrac{\beta_{t}}{\alpha_{t}-1}$.
By examining equations (3) and (4), we obtain:
$\begin{aligned} \left|\mathbb{E}\left[y_{t}\right]-y_{\star}\right|&=\left(-\ln(\frac{h}{p+h})\right)^{1/k}\left|\mathbb{E}\left[\left(\frac{1}{\theta_{t}}\right)^{1/k}\right]-\left(\frac{1}{\theta_{\star}}\right)^{1/k}\right|\\ &=\left(-\ln(\frac{h}{p+h})\right)^{1/k}\left|\mathbb{E}\left[\left(\frac{1}{\theta_{t}}\right)^{1/k}-\left(\frac{1}{\theta_{\star}}\right)^{1/k}\right]\right|\\ &\leq\left(-\ln(\frac{h}{p+h})\right)^{1/k}\mathbb{E}\left[\left|\left(\frac{1}{\theta_{t}}\right)^{1/k}-\left(\frac{1}{\theta_{\star}}\right)^{1/k}\right|\right]\\ &\leq\left(-\ln(\frac{h}{p+h})\right)^{1/k}\cdot\frac{1}{k}\cdot\min\left\{L,\frac{1}{\theta_{*}}\right\}^{1/k-1}\cdot\mathbb{E}\left[\left|\left(\frac{1}{\theta_{t}}\right)-\left(\frac{1}{\theta_{\star}}\right)\right|\right]\end{aligned}$ (5)
The first inequality comes from jensen inequality $|\mathbb{E}[X]|\leq\mathbb{E}[|X|]$. The second inequality comes from Lipschitz inequality and $y_{t}$’s lower bound stated earlier in Lemma 1.
Lemma 4.
For each $t\in[T]$, with probability $\geq 1-\delta/T$, the realization of demand $D_{t}\sim\operatorname{Weibull}(\theta_{\star})$ lies in
$\underline{D}=\left(\tfrac{\ln{\left(\tfrac{2T}{2T-\delta}\right)}}{\theta_{\star}}\right)^{\frac{1}{k}},\qquad\overline{D}=\left(\tfrac{\ln{\left(\tfrac{2T}{\delta}\right)}}{\theta_{\star}}\right)^{\frac{1}{k}}.$
Lemma 4 is proved in Appendix A.2. This lemma ensures that, with high probability, the demand realizations are confined within the specified range, which is crucial for later analysis. Next, we provide confidence bound for how close the $\frac{1}{\theta_{t}}$ and its mean $\frac{\beta_{t}}{\alpha_{t}-1}$ is. Ideally, as $t$ increases, $\theta_{t}$ will converge to $\theta_{*}$ and $\frac{1}{\theta_{t}}$ will converge to $\frac{\beta_{t}}{\alpha_{t}-1}$. The following lemma shows the rate of convergence as follows:
Lemma 5.
For any $t\in[T]$ and for any realization of $\theta_{\star}$,
$\mathbb{P}\left(\left|\tfrac{\beta_{t}}{\alpha_{t}-1}-\tfrac{1}{\theta_{\star}}\right|\geq\sqrt{\ln{\left(\tfrac{2t^{2}}{\delta}\right)}}\left(\overline{D}^{k}+\tfrac{2}{\theta_{\star}}\right)\sqrt{\tfrac{t}{\left(\alpha_{t}-1\right)^{2}}}\right)\leq\tfrac{\delta}{t^{2}}.$
Lemma 5 is proved in Appendix A.3. This lemma provides a probabilistic bound on the estimation error of $\frac{1}{\theta_{t}}$, which is key in assessing the accuracy of the order quantity decisions over time. Combining Lemmas 4 and 5, we bound
$\left|\mathbb{E}\left[y_{t}\right]-y_{\star}\right|$ $\leq\left(-\ln(\frac{h}{p+h})\right)^{1/k}\cdot\frac{1}{k}\cdot\min\left\{L,\frac{1}{\theta_{*}}\right\}^{1/k-1}\sqrt{\ln{\left(\frac{2t^{2}}{\delta}\right)}}\left(\overline{D}^{k}+\frac{2}{\theta_{\star}}\right)\sqrt{\frac{t}{\left(\alpha_{t}-1\right)^{2}}}$
$\leq\left(-\ln(\frac{h}{p+h})\right)^{1/k}\left(\overline{D}^{k}+\frac{2}{\theta_{\star}}\right)\cdot\frac{1}{k}\cdot\min\left\{L,\frac{1}{\theta_{*}}\right\}^{1/k-1}\sqrt{2\ln{\left(\frac{T}{\delta}\right)}}\sqrt{\frac{t}{\left(\alpha_{t}-1\right)^{2}}}.$
Lemma 6 (Chuang and Kim (2023)).
$\alpha_{t}$ $=\alpha_{0}+\sum_{i=0}^{t-1}\delta_{i}=\alpha_{0}+\sum_{i=0}^{t-1}\mathbb{E}_{\theta_{\star}}\left[\delta_{i}\mid H_{i-1}\right]+\sum_{i=0}^{t-1}\left(\delta_{i}-\mathbb{E}_{\theta_{\star}}\left[\delta_{i}\mid H_{i-1}\right]\right)=\alpha_{0}+\sum_{i=0}^{t-1}\left(1-e^{-\theta_{\star}y_{i}^{k}}\right)+M_{t}$
where $M_{t}=\sum_{i=0}^{t-1}\left(\delta_{i}-\mathbb{E}\left[\delta_{i}\mid\mathcal{H}_{i-1}\right]\right)-1$.
From the above lemma 6, we can see that as long as $y_{t}$ has a lower bound, we are able to derive the upper bound for regret. Then, we plug back the lower bound to the definition of $\alpha_{t}$ in Lemma 6 to analyze the term $\left|\alpha_{t}-1\right|$, which analysis is referred in Appendix A.9. To establish the regret, we use the technique of truncating $T$ as follows. Denote $T_{0}=64\left(1-\exp\{-\theta_{\star}L^{k}\}\right)^{-2}\ln{\left(\frac{T}{\delta}\right)}.$ According to the proof of Lemma 11 in the Appendix , we have $\forall t\in[T],$
$\frac{\sqrt{t}}{\alpha_{t}-1}\leq\begin{cases}\frac{\sqrt{T_{0}}}{\alpha_{0}-1}&t\leq T_{0},\\ \frac{\sqrt{t}}{\frac{1}{2}t(1-\exp{\{-\theta_{\star}L^{k}\}})}=\frac{2}{\sqrt{t}\cdot(1-\exp{\{-\theta_{\star}L^{k}\}})}&t>T_{0}.\end{cases}$
Therefore,
$\sum_{t=1}^{T}\frac{\sqrt{t}}{\alpha_{t}-1}=\sum_{t=1}^{T_{0}}\frac{\sqrt{t}}{\alpha_{t}-1}+\sum_{t=T_{0}}^{T}\frac{\sqrt{t}}{\alpha_{t}-1}\\ \leq T_{0}\cdot\frac{\sqrt{T_{0}}}{\alpha_{0}-1}+\frac{2}{(1-\exp{\{-\theta_{\star}L^{k}\}})}\cdot\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\\ \leq T_{0}^{\frac{3}{2}}\cdot\frac{1}{\alpha_{0}-1}+4\left(1-\exp\{-\theta_{\star}L^{k}\}\right)^{-1}\sqrt{T}$
3.2.1 Key Step 3: Regret Decomposition via Lipschitz Continuity